24th September 2008 APM partially new tables 11:30-13:00 Present: MJI, JRL, PSB, MR, DWE, EGS, RGM Apologies: NAW, STH Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Telecon with JAC 4. ESO surveys workshop 5. VOTC UK workshop 6. AOB Minutes ------- 1. Actions from last meeting STH has not finished checking that the Plone astrometry technical pages are correct - ongoing <<<< (see minutes from 080332 for further details) STH has not included a graphic showing the size of the photometric radial distortion correction in the Plone technical pages nor checked that the rest of the photometry pages are up to date <<<< JRL has not produced a web page of HawkI publicity for CASU - ongoing <<<< MR has installed JRL's script to flag summit-rejected MSBs during raw WFCAM data ingestion before WFCAM goes back on MJI has discussed topics and suitable time/date for a telecon with JAC (see item 3) MJI have not finished off the stage-I software release pages - ongoing <<<< EGS EGS has written the presentation for the SADT+P2PP meeting and rammed home the appropriate points PSB helped STH prepare for his VISTA Paranal trip, but did not have anything to do with his laundry JRL have yet to attend the Cambridge VOTC meeting since it hasn't happened MR yet - ongoing <<<< STH has not made the photometric calibration paper available on the Plone and VDFS sites - ongoing <<<< MJI has not finished off the technical paper - ongoing <<<< JRL investigated object masking sky strategy using 20080521 as a test night MJI has not investigated the WFCAM bad pixel map issue any further - ongoing <<<< PSB has checked that the new machines' power consumption and fan speed are regulateable. Cooling occurs as it is needed. If you exercise a machine, then its fan whizzes around faster 2. Comments on WFAU minutes 3. Telecon with JAC 4. ESO surveys workshop The directory structure proposed by ESO for the transfer of data was thought reasonable. MJI will email the appropriate person to inform them. <<<< We would like the first few nights of SV data as soon as possible so that we can provide prompt feedback. MJI will contact the appropriate person. <<<< 2 disks were mentioned - presumably we would get a partially filled disk early. PSB will send some of the already processed VIRCAM test data to Edinburgh so that they can test ingestion. <<<< There was a discussion about hard-coding Ndit=1 into the template. JRL to contact Steven Beard? There probably should be an action about this. <<<< There was also some additional stuff that needed to be added to STH's observing wish-list. MJI to coordinate this? <<<< 5. VOTC UK workshop 6. AOB Continuing actions ------------------ STH check that Plone astrometry technical pages are correct STH include graphic showing the size of the photometric radial distortion correction in the Plone technical pages and check rest of photometry pages are up to date JRL produce web page of HawkI publicity for CASU MJI finish off stage-I software release pages EGS JRL attend Cambridge VOTC meeting on 25-26 September MR STH make photometric calibration paper available on Plone and VDFS sites MJI finish off technical paper MJI investigate WFCAM bad pixel map issue further New actions ----------- MJI contact ESO about CASU accepting proposed directory structure for transferring data MJI contact ESO about acquiring first few nights of SV data as soon as possible after it is taken PSB send some processed VIRCAM test data to Edinburgh JRL contact Steven Beard about hard-coding Ndit=1 into template ??? MJI to coordinate adding items to STH's observing wish-list